OpenStudy (anonymous):

Find the product? (3p-7)(3p+7) A.9^2+49 B.9^2-49 C.9^2-42P-49 D.9^2+42P+49

OpenStudy (anonymous):

you would have to know that \[(a-b)(a+b)=a^2-b^2\]

OpenStudy (anonymous):

9p^2+21p-21p-49=9p^2-49

OpenStudy (anonymous):

using that we have \[(3p-7)(3p+7)=(3p)^2-(7)^2\] \[=9p^2-49\]

OpenStudy (anonymous):

thancks Beautiful Girls:)

OpenStudy (anonymous):

yw